# Knowledge-Guided Prompt Learning for Request Quality Assurance in Public Code Review

### v1

@@ 4.2.3. Knowledge-guided Prompt
In order to optimize the code representation and better guide the knowledge embedded in the request text and code snippets, we use the method of continuous prompting (soft prompt) (Liu et al., 2023b) to introduce external knowledge. Specifically, we first extract the professional term from the request using string matching, then initialize the $P_{prefix}$ vector prefix by concatenating the knowledge interpretation representation, and concatenate the representation vector of the data flow graph of the code snippet after this prefix vector. Finally, the whole vector is represented as the model code snippet instead of the original code snippet. Further details on these extensions can be found in Section 4.3.
@@ 4.3.1. Knowledge-Guided Prompt Learning
Soft prompting is one of the prompting paradigms, which operates in the embedding space of the model and adjusts the prompt parameters for downstream tasks. Currently, this method has demonstrated superior performance on other non-natural language data, such as image encoding and visual question answering (Liu et al., 2023a; Tsimpoukelli et al., 2021). This indicates that soft prompts are well-suited for learning and optimizing information with such graph structures, owing to the ability of vector embeddings in the embedding space to better adapt to non-natural language structures.
Since the code snippets and text in public code review requests contain a large amount of domain-specific knowledge, which is not fully learned by traditional language models, this study uses the labels of professional terms as a benchmark. According to the vocabulary of pre-trained language models, words not in the vocabulary are first selected, and their professional knowledge definitions are obtained from Wikipedia. These words are then saved as a separate external knowledge file. When processing individual request bodies, since tokenization may split words not present in the language model into roots and suffixes, words are split based on spaces using string manipulation. Furthermore, the split words are subjected to fuzzy casing, tense, and part-of-speech transformations. Next, the processed text string is used to extract several domain-specific knowledge words contained in each request body through string matching. Finally, the definitions of domain-specific knowledge words are concatenated and their representations are obtained, which is then fed into the model in the form of prefix vectors to assist in learning the knowledge in the request body.
@@ 1. Natural language text $text_{i}$, composed of the title and description (text part) of the request;
1:$req=\left(title,\ text,\ code\right)$
2:$r_{\widetilde{prompt}}$
3:$Graph_{c}\leftarrow Graph\left(code\right)$, $Graph\left(code\right)$ represents the data flow graph of the code snippet;
4:$h_{Graph_{c}}\leftarrow PLM\left(Graph_{c}\right)$, $h_{Graph_{c}}$ represents the code hidden vector obtained through the encoding model;
5:$knowledge=req\cap\text{wiki}$;
6:$P_{prefix}=PLM\left(knowledge\right),\ PLM\left(knowledge\right)$ represents the representation vector obtained through the encoding model;
7:$r_{\widetilde{prompt}}=\left\{T\left(\cdot\right),\ \widetilde{title},\ \widetilde{text},\ P_{prefix},\ h_{Graph_{c}}\right\}$;
@@
The method for processing the code snippet and constructing the prefix vector is shown in Algorithm 1. In Step ①, a code snippet $code_{i}$ is parsed into a data flow graph to represent the dependency relationship among code variables (as shown on the right side of Fig. 5). Then, in Step ②, the variable relationship graph is input into a pre-trained code encoding model for code tasks to obtain the hidden vector of the code snippet. In this work, the GraphCodeBERT (Guo et al., 2021) model is used as the encoding model. In Step ③, professional knowledge terms are extracted from the request using string matching. Step ④ defines a prefix vector $P_{prefix}$, then initializes the prefix vector $P_{prefix}$ with the representation vector of knowledge processed by the encoding model. Finally, the entire vector replaces the original code snippet as the model’s code representation.
@@
Here, $P_{idx}$ is used to represent the sequence of prefix indices, and $\sigma$ is used to represent the non-linear output obtained from the hidden layers of the code pre-trained language model. To address the issue of low performance and high variance caused by the random initialization of $P_{prefix}$, we initialize the prefix vector $P_{prefix}$ with the text of external knowledge and then use an immutable data flow graph based on semantic relationships to represent the meaning of individual code snippets. This allows us to guide the model to learn professional knowledge from the request text through external knowledge and to globally and locally understand the meaning of each code snippet through the form of prefix vectors.
@@
In order to measure the efficiency of Algorithm 1, we analyze its time complexity. To this end, we consider the computational overhead in each step. In Step ①, the size of a code snippet determines the time complexity to convert code into a data flow diagram. If there are $n$ variables or code fragments in the code snippet, generating a dependency diagram requires checking the dependencies between each pair of variables, and consequently the time complexity is $O\left(n^{2}\right)$. Then, in Step ②, the time complexity of GraphCodeBERT generally depends on the architecture of its pre-trained model and the input graph size of a data flow diagram. Overall, the complexity of Step 2 can be seen as $O\left(V+E\right)$, in which $V$ is the number of nodes in the graph, and $E$ is the number of edges. The complexity of this step is also $O\left(n^{2}\right)$, because the encoding process is proportional to the size of the graph. Step ③ uses string matching whose time complexity is $O\left(s\right)$, in which $s$ is the length of the request text. The time complexity of Step ④ depends on the size of the encoding model and the length of the knowledge representation vector. The process of initializing a vector can be regarded as a constant time operation, and the complexity is $O\left(d\right)$, in which $d$ is the length of the vector. Therefore, the overall time complexity of the algorithm can be expressed as $O\left(n^{2}+s+d\right)$. Since the complexity of language model representation is much higher than that of string matching, i.e., $O\left(n^{2}\right)\gg O\left(s\right)$ and $O\left(n^{2}\right)\gg O\left(d\right)$, and the time complexity can also be expressed as $O\left(n^{2}\right)$.
@@
Compared to our preliminary study (Chen et al., 2024) where the time complexity is also $O\left(n^{2}\right)$, it can be found that according to the experimental results in Section 6, our KP-PCR method can improve the task accuracy of request necessity prediction and tag recommendation.
@@ 4.3.2. Loss Function
Figure 6. Trainable and Frozen Parameters
@@
where $\varphi$ is the predicted probability distribution, and $\varphi_{j}\left(label_{r}\right)$ represents the mapping from the $j$-th $h_{MASK}$ token to the requested label word $label$ in the input sequence. $\phi$ denotes the parameters of the language model that are updatable. $n$ is the number of masked positions in the descriptive prompt template $T\left(req\right)$.
@@ 5.1. Research Questions
We randomly extract 20 request bodies and their corresponding task tags from the existing public code review data. GPT-3.5 and GPT-4.0 are required to select tags according to the task definition in the absence of external resources and with a few references. We explore how GPT-3.5 and GPT-4.0 (1) perform under different prompt descriptions and (2) perform common code review quality assurance tasks by changing the prompt description of the task.
@@ 5.2.1. Dataset Description
To evaluate the effectiveness of our KP-PCR, a dataset on public code review from Stack Exchange is selected, a reputable well-known software information platform. The selection of the dataset follows the criteria of being open, authoritative, and current. Stack Exchange is a large network of software question and answer (Q&A) websites covering various domains, such as programming, Linux & Unix, and code review. Each website within this network focuses on a specific topic, and the questions, answers, and users are subject to reputation-based recognition within their respective communities.
In the field of software engineering, there are other software engineering information platforms such as Stack Overflow and WordPress. As we are dedicated to providing intelligent service methods for the code review process, the most popular public code review website on Stack Exchange, Code Review [^1] https://codereview.stackexchange.com/, is used as the data source for this study. The dataset from Code Review used for this study is the latest and largest original data version available from 2011 to 2023222https://archive.org/download/stackexchange/codereview.stackexchange.com.7z. It records the main content and quality label of PCR requests in separate XML files.
Fig. 7 introduces the topics of the Code Review website, which is positioned as a peer code review conducted in the form of SQA (Software Quality Assurance). It aims to facilitate a quick and convenient code review process to accelerate the exchange of professional knowledge among peer developers. Any developer can initiate a review request in the community, and peers will assess the necessity of the review based on their expertise, presenting it in the form of contribution scores on each request.
Figure 7. Code Review Website Introduction.
[…]
which is positioned as a peer code review conducted in the form of SQA
@@ 5.2.2. Data Preprocessing
Following previous research, we have utilized the common ten-fold cross-validation method to split the dataset. Specifically, we initially randomly have divided the dataset into 10 equally sized subsets. One of these subsets is selected as the test set. The remaining data samples are split into training and validation sets in an 8:2 ratio. This process is repeated during model training to test the model’s effectiveness, and the results is averaged to assess model performance. In summary, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. Detailed statistical information is provided in Table 2.
@@ 5.2.3. Evaluation Metrics
(1) $Top-k$ Precision $Precision@k$: This metric is commonly used in recommendation-related methods, measuring the proportion of true labels among the top $k$ labels recommended by a model. For a single request, the calculation of its $Top-k$ Accuracy is shown in Equation 4 which computes $Precision@k_{i}$ by calculating the intersection of predicted labels $vk_{i}$ with the corresponding true labels $Ttag_{i}$ for each request object. Since a public code review dataset contains multiple requests, when calculating accuracy metrics for the entire dataset, the average accuracy results for all requests need to be taken into account. The specific calculation method is shown in Equation 5 which represents the average accuracy of all request objects, calculated by first computing $Precision@k_{i}$ for each request object, and then averaging them to represent $Precision@k$.
@@
(2) $Top-k$ Recall $Recall@k$: This metric is used to evaluate the model’s ability to predict the complete set of true labels. Specifically, it measures the proportion of true labels that are contained in the predicted labels. The $Top-k$ Recall for a single request object, $Recall@k_{i}$, is calculated as shown in Equation 6. In this work, due to the true label length range being 3-5, and following the metrics settings of recommendation-related methods, $k$ is set to 3, 5, and 10 in this study. Similarly, since there are multiple request objects in the public code review dataset, the performance calculation of Recall metric $Recall@k$ for the entire dataset is shown in Equation 7. The average recall rate of all request objects is calculated to represent the model’s performance on this metric.
(6) $Recall@k_{i}=\left\{\begin{array}[]{l}\left|\frac{Ttag_{i}\cap vk_{i}}{k}\right|,if\left|Tag_{i}\right|<k,\\ \left|\frac{Ttag_{i}\cap vk_{i}}{Ttag_{i}}\right|,otherwise.\\ \end{array}\right.$
@@
(3) $Top-k$ Macro-Average $F1@k$: This metric is commonly used to evaluate a model’s overall performance on precision and recall by calculating the weighted average of these two metrics. It provides a balanced view of a model’s performance by considering both precision and recall. As described in Equation 8, the calculation method for the macro-average $F1@k_{i}$ of a single request object is similar to the previous two metrics, where $k$ is set to 3, 5, and 10 in this study. Similarly, for the overall macro-average metric $F1@k$ of the public code review dataset, its calculation is shown in Equation 9, derived from the average $F1$ value of all request objects.
@@ 5.3.2. Baselines
Our previous study is also listed for comparison. The fine-tuning parameters provided by the baseline methods are used, and their optimal values are adopted in experimenting baselines. For baseline methods which source code was not provided, this study reproduced the model design based on Pytorch framework. The related experiments in this study are conducted on four NVIDIA Titan XP GPUs. All hardware and software environments are shown in Table 4.
@@ 6.1. Request Necessity Prediction Subtask (RQ1)
(1) Compared to baseline methods, KP-PCR shows improved performance in the review necessity prediction subtask. As shown in Table 5, KP-PCR achieves a 0.7%-8.3% performance improvement over the second-best method (UniPCR) by introducing the code snippet optimization module. This demonstrates the effectiveness of our KP-PCR. By learning the representation of code snippets within the overall content through soft prompts in the prefix vectors, KP-PCR understands the differences in various code snippets under the review necessity criteria. During prediction, the representation in the learned code snippet prefix vectors enhances the prediction performance for review necessity, resulting in KP-PCR’s improved performance in this subtask.

### v3

@@ 4.2.3. Knowledge-Guided Prompt
In order to optimize the code representation and better guide the knowledge embedded in the request text and code snippets, we use the continuous prompting method (soft prompt) (Liu et al., 2023b) to introduce knowledge guidance. Specifically, we first extract the professional term from the request using string matching, then initialize the $P_{prefix}$ vector prefix by concatenating the knowledge interpretation representation and concatenate the representation vector of the program dependence graph of the code snippet after this prefix vector. Finally, the whole vector is represented as the model code snippet instead of the original code snippet. Further details on these extensions can be found in Section 4.3.
@@ 4.3.1. Knowledge-Guided Prompt Learning
Soft prompting is one of the prompting paradigms, which operates in the embedding space of the model and adjusts the prompt parameters for downstream tasks. Currently, this method has shown superior performance in other data not in natural language, such as image encoding and visual question answering (Liu et al., 2023a; Tsimpoukelli et al., 2021). This indicates that soft prompts are well-suited for learning and optimizing information with such graph structures, owing to the ability of vector embeddings in the embedding space to better adapt to non-natural language structures. The PCR task is not merely about code quality detection or generation; it involves a multi-step code review process, which includes predicting whether the code needs review and how to select appropriate labels for the code review task. Knowledge-guided learning in PCR can help the model understand the context of the code and identify common algorithms, functions, or libraries involved in the code snippets, thereby enhancing the model’s comprehension and judgment of code review requests.
To enhance the model’s performance by integrating external knowledge, our framework uses LLM, specifically introducing the DeepSeek-7B. The reason for choosing this model is that it has demonstrated proven adaptability in code understanding tasks, and its parameter scale is relatively small, requiring less computational resources. It can be deployed in an environment with at least 16GB of GPU memory. By performing label retrieval on Wikipedia for the dataset, we have obtained 1,000 articles related to software engineering, programming languages, algorithms, and technical concepts. These articles are then preprocessed to ensure that their format was suitable for the LLM’s input requirements. The data was cleaned to remove irrelevant content, leaving only technical definitions, code examples, and domain-specific knowledge. DeepSeek-7B is then fine-tuned using a supervised learning approach, with a learning rate of $2\times 10^{-5}$, a batch size of 8, and trained for 10 epochs. Regarding the consumption of inference time, after testing under the condition of a 512-token input, the response time for every 1000 rounds is 8 minutes, and the average response time per round is 480 milliseconds. The additional computational overhead brought by using LLM is reasonable and controllable.
This study uses the labels of specialized terminology as a benchmark and selects words that are not present in the vocabulary of the pre-trained language model. First, string operations are used to split words based on spaces. Additionally, the split words undergo fuzzy case, tense, and part-of-speech transformations. Next, the processed text strings are used for string matching to extract domain-specific knowledge words contained in each request body, which are then input into the fine-tuned DeepSeek model to obtain their expert definitions. Finally, the definitions of domain-specific knowledge words are concatenated to form their representations as input into the model. And they work in the form of prefix vectors to aid in learning the knowledge within the request body.
@@ 1. Natural language text $text_{i}$, composed of the title and description (text part) of the request;
[…]
@@
The method for processing the code snippet and constructing the prefix vector is shown in Algorithm 1. In Step ①, a code snippet $code_{i}$ is parsed into a program dependence graph to represent the dependency relationship among code variables (as shown on the right side of Fig. LABEL:fig:dataflow). Then, in Step ②, the variable relationship graph is input into a pre-trained code encoding model for code tasks to obtain the hidden vector of the code snippet. In this work, the GraphCodeBERT (Guo et al., 2021) model is used as the encoding model. In Step ③, professional knowledge terms are extracted from the request using string matching. Step ④ defines a prefix vector $P_{prefix}$, then initializes the prefix vector $P_{prefix}$ with the representation vector of knowledge processed by the encoding model. Finally, the entire vector replaces the original code snippet as the model’s code representation.
@@
Here, $P_{idx}$ is used to represent the sequence of prefix indices, and $\sigma$ is used to represent the non-linear output obtained from the hidden layers of the code pre-trained language model. To address the issue of low performance and high variance caused by the random initialization of $P_{prefix}$, we initialize the prefix vector $P_{prefix}$ with the text of knowledge guidance and then use an immutable program dependence graph based on semantic relationships to represent the meaning of individual code snippets. This allows us to guide the model to learn professional knowledge from the request text through knowledge guidance and to globally and locally understand the meaning of each code snippet through the form of prefix vectors.
@@
In order to measure the efficiency of Algorithm 1, we analyze its time complexity. To this end, we consider the computational overhead in each step. In Step ①, the size of a code snippet determines the time complexity to convert code into a program dependence graph. If there are $n$ variables or code fragments in the code snippet, generating a program dependence graph requires checking the dependencies between each pair of variables, and consequently the time complexity is $O\left(n^{2}\right)$. Then, in Step ②, the time complexity of GraphCodeBERT generally depends on the architecture of its pre-trained model and the input graph size of a program dependence graph. Overall, the complexity of Step 2 can be seen as $O\left(V+E\right)$, in which $V$ is the number of nodes in the graph, and $E$ is the number of edges. The complexity of this step is also $O\left(n^{2}\right)$, because the encoding process is proportional to the size of the graph. Step ③ uses string matching whose time complexity is $O\left(s\right)$, in which $s$ is the length of the request text. The time complexity of Step ④ depends on the size of the encoding model and the length of the knowledge representation vector. The process of initializing a vector can be regarded as a constant time operation, and the complexity is $O\left(d\right)$, in which $d$ is the length of the vector. Therefore, the overall time complexity of the algorithm can be expressed as $O\left(n^{2}+s+d\right)$. Since the complexity of language model representation is much higher than that of string matching, i.e., $O\left(n^{2}\right)\gg O\left(s\right)$ and $O\left(n^{2}\right)\gg O\left(d\right)$, and the time complexity can also be expressed as $O\left(n^{2}\right)$.
@@
Our preliminary study (Chen et al., 2024), where the time complexity is also $O\left(n^{2}\right)$, has an actual runtime of approximately 8 hours per epoch. The actual runtime of each epoch for our KP-PCR method is about 8 hours and 20 minutes, which is nearly the same. However, it can be found that according to the experimental results in Section 6, our KP-PCR method can improve the task accuracy of request necessity prediction and tag recommendation.
@@ 4.3.2. Loss Function
Figure 5. Trainable and Frozen Parameters.
@@
where $\varphi$ is the predicted probability distribution, and $\varphi_{j}\left(label_{r}\right)$ represents the mapping from the $j$-th $h_{MASK}$ token to the requested label word labellabel in the input sequence. $\phi$ denotes the parameters of the language model that are updatable. $n$ is the number of masked positions in the descriptive prompt template $T\left(req\right)$.
@@ 5.1. Research Questions
we fine-tuned Llama3, CodeLlama and DeepSeek using the entire training dataset, and then evaluated them on the test set, comparing their performance with our model.
@@ 5.2.1. Dataset Description
To evaluate the effectiveness of our KP-PCR, experimental data is from the Code Review site within Stack Exchange, a reputable and well-known information-sharing platform. The selection of the dataset follows the criteria of being open, authoritative, and current. Stack Exchange is a large network of question and answer (Q&A) websites covering a wide range of topics, including but not limited to software domains such as programming, Linux & Unix, and code review. Each website within this network focuses on a specific theme, and the questions, answers, and users are subject to reputation-based recognition within their respective communities.
In the field of software engineering, there are several traditional public code review platforms such as GitHub and Reddit. Stack Exchange is a Q&A-based community that encourages quick feedback. Any developer can submit code snippets without the need to provide a complete project or engage in complex version control, making it highly accessible and ideal for extracting smaller, isolated code review requests. This contrasts with GitHub, which is more focused on project-driven collaboration, and Reddit, where the review process is often informal and may lack in-depth technical feedback. Unlike Reddit, where feedback is typically provided through informal post comments and may lack the necessary technical depth, Stack Exchange provides a more structured environment for feedback. Users are encouraged to give specific, detailed responses, and reviews often involve technical explanations and recommendations. This level of technical detail makes Stack Exchange an ideal environment for extracting meaningful data for public code review tasks.
As we are dedicated to providing intelligent service methods for the code review process, the most popular public code review website on Stack Exchange, Code Review [^2] https://codereview.stackexchange.com/, is used as the data source for this study. The dataset from Code Review used for this study is the latest and largest original data version available from 2011 to 2023333https://archive.org/download/stackexchange/codereview.stackexchange.com.7z. It records the main content and quality label of PCR requests in separate XML files.
Code Review website, which is positioned as peer code review conducted in the form of SQA (Software Quality Assurance). It aims to facilitate a quick and convenient code review process to accelerate the exchange of professional knowledge among peer developers. Any developer can initiate a review request within the community. Professional reviewers in the community then browse through and assess the necessity of the request. Subsequently, the developer selects appropriate tags for the request to match the corresponding reviewers. The dataset collects information on the requests content, the necessity of the request, and the corresponding tags.
@@ 5.2.2. Data Preprocessing
(3) Necessity Determination: The distribution of request vote scores on the Code Review site is analyzed. It is observed that the distribution follows a power-law pattern, exhibiting a “long-tail” characteristic in which most posts receive relatively low scores, while a small proportion obtain significantly higher scores. In such distributions, the upper percentile (e.g., top 40–50%) of items are often regarded as high-quality “core contributions” recognized by the community (Newman, 2005). The analysis showed that, in the Code Review community, requests with a score of at least 4 corresponded precisely to the top 43%. Therefore, requests with a vote score of at least 4 are classified as necessary requests, whereas those below this threshold are considered unnecessary requests. This resulted in a total of 26,632 necessary requests and 35,056 unnecessary requests.
@@
Following previous research, we have utilized the common ten-fold cross-validation method to split the dataset. Specifically, we initially randomly have divided the dataset into 10 equally sized subsets. One of these subsets is selected as the test set. The remaining data samples are split into training and validation sets in an 8:1 ratio. This process is repeated during model training to test the model’s effectiveness, and the results is averaged to assess model performance. In summary, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. Detailed statistical information is provided in Table 2.
@@ 5.2.3. Evaluation Metrics
(1) TopkTop-k Precision $Precision@k$: This metric is commonly used in recommendation-related methods, measuring the proportion of true labels among the top $k$ labels recommended by a model. For a single request, the calculation of its TopkTop-k Accuracy is shown in Equation 4 which computes $Precision@k_{i}$ by calculating the intersection of predicted labels $vk_{i}$ with the corresponding true labels $Ttag_{i}$ for each request object. Since a public code review dataset contains multiple requests, when calculating accuracy metrics for the entire dataset, the average accuracy results for all requests need to be taken into account. The specific calculation method is shown in Equation 5 which represents the average accuracy of all request objects, calculated by first computing $Precision@k_{i}$ for each request object, and then averaging them to represent $Precision@k$.
@@
(2) TopkTop-k Recall $Recall@k$: This metric is used to evaluate the model’s ability to predict the complete set of true labels. Specifically, it measures the proportion of true labels that are contained in the predicted labels. The TopkTop-k Recall for a single request object, $Recall@k_{i}$, is calculated as shown in Equation 6. In this work, due to the true label length range being 3-5, and following the metrics settings of recommendation-related methods, $k$ is set to 3, 5, and 10 in this study. Similarly, since there are multiple request objects in the public code review dataset, the performance calculation of Recall metric $Recall@k$ for the entire dataset is shown in Equation 7. The average recall rate of all request objects is calculated to represent the model’s performance on this metric.
[…]
@@
(3) TopkTop-k Macro-Average $F1@k$: This metric is commonly used to evaluate a model’s overall performance on precision and recall by calculating the weighted average of these two metrics. It provides a balanced view of a model’s performance by considering both precision and recall. As described in Equation 8, the calculation method for the macro-average $F1@k_{i}$ of a single request object is similar to the previous two metrics, where $k$ is set to 3, 5, and 10 in this study. Similarly, for the overall macro-average metric $F1@k$ of the public code review dataset, its calculation is shown in Equation 9, derived from the average $F1$ value of all request objects.
@@ 5.3.2. Baselines
Our previous study is also listed for comparison. In this paper, the source code provided by the baselines is used to fine-tune the parameters and obtain the optimal values to represent the experimental results of the baselines. For baselines which source code was not provided, this study reproduced the model design based on Pytorch framework. The related experiments in this study are conducted on four NVIDIA Titan XP GPUs. All hardware and software environments are shown in Table 4.
@@ 6.1. Request Necessity Prediction Subtask (RQ1)
(1) Compared to baseline methods, KP-PCR shows improved performance in the review necessity prediction subtask. As shown in Table 5, KP-PCR achieves a 2.3%-8.4% performance improvement over the second-best method (UniPCR) by introducing the code snippet optimization module. This demonstrates the effectiveness of our KP-PCR. By learning the representation of code snippets within the overall content through soft prompts in the prefix vectors, KP-PCR understands the differences in various code snippets under the review necessity criteria. During prediction, the representation in the learned code snippet prefix vectors enhances the prediction performance for review necessity, resulting in KP-PCR’s improved performance in this subtask.